If Car A leaves at 1pm traveling 50 mph and Car B leaves at 2 pm traveling 65 mph, what time will Car B pass Car A?
(Please show the work.)

Answer by stanbon(75887)

If Car A leaves at 1pm traveling 50 mph and Car B leaves at 2 pm traveling 65 mph, what time will Car B pass Car A?
--------------
Car A DATA:
rate = 50 mph ; time = x hrs ; distance = 50x miles
-------------
Car B DATA:
rate = 65 mph ; time = (x-1) hrs ; distance = 65(x-1) miles
-------------
Equation:
distance = distance
50x = 65(x-1)
60x = 65x - 65
5x = 65
x = 13 hrs
-----------------
Time when they have equal distance:
1 pm + 13 hrs = 2 am
=========================
Cheers,
Stan H.